# Recurrent Drafter for Fast Speculative Decoding in Large Language Models

### v4

@@ Abstract
We present Recurrent Drafter (ReDrafter), an advanced speculative decoding approach that achieves state-of-the-art speedup for large language models (LLMs) inference. The performance gains are driven by three key aspects: (1) leveraging a recurrent neural network (RNN) as the draft model conditioning on LLM’s hidden states, (2) applying a dynamic tree attention algorithm over beam search results to eliminate duplicated prefixes in candidate sequences, and (3) training through knowledge distillation from the LLM. ReDrafter accelerates Vicuna inference in MT-Bench by up to 3.5x with a PyTorch implementation on Nvidia H100 GPUs. To demonstrate its practicality in production environments, we integrate ReDrafter into TensorRT-LLM, reaching up to 2.5x speedup on H100 GPUs. We also validated its effectiveness for on-device applications by implementing the approach in MLX and benchmarking performance on Metal GPUs in Apple Silicon chips, achieving up to 2.3x speedup. We summarize our experimental results in Figure 1.
@@ 1 Introduction
Figure 1: Inference speedup on MT-bench for Vicuna models are shown for three ReDrafter implementations: (left) PyTorch on Nvidia H100 GPU, (mid) TensorRT-LLM on Nvidia H100 GPU, (right) and MLX on Apple’s M2 Ultra Metal GPU. We compare with EAGLE (Li et al., 2024a), Medusa (Cai et al., 2024), and auto-regressive in PyTorch implementations.
@@
In a PyTorch implementation, ReDrafter accelerates Vicuna inference by up to 3.5x compared to the autoregressive method in MT-Bench on Nvidia H100 GPUs, achieving state-of-the-art performance. Additionally, we demonstrate the effectiveness of ReDrafter in production environments through two key use cases. The first is a production-level deployment on TensorRT-LLM, designed to manage high traffic with long context lengths. In this implementation, tensor parallelism and continuous batching are utilized to ensure the system efficiently handles larger volume of requests while maintaining low latency. In this scenario, ReDrafter achieves up to 2.5x speedup on the MT-bench benchmark. The second use case focuses on an on-device approach using MLX on Metal GPUs within Apple Silicon chips. Despite the limited compute resources in this setup, we observed a memory bottleneck. ReDrafter effectively mitigates this bottleneck, resulting in up to 2.3x speedup, demonstrating its capability to optimize performance in resource-constrained environments.
@@ 2 Related Work on Speculative Decoding
Another thread of approaches employs a unified strategy by attaching the draft model to the LLM, making them dependent (Stern et al., 2018; Santilli et al., 2023; Cai et al., 2024). This is an efficient design choice when the draft model is not intended for standalone use, allowing it to leverage additional computational resources for local prediction by conditioning on the LLM. Among these methods, Cai et al. (2024) proposes to use $T$ independent draft heads to predict next $T$ tokens, utilizing GPU’s extra parallel computing power. This excessive computational effort may not yield proportional speedups, as independent predictions become less accurate as $T$ increases, resulting in suboptimal predictive accuracy and a lower acceptance rate from the LLM. An alternative approach is to enhance prediction accuracy by incorporating recurrence into the draft model to capture local dependencies (Bhendawade et al., 2024; Li et al., 2024a; b). However, the reduced parallelism due to recurrence leads to lower GPU utilization, introducing overhead that diminishes speedup gains, even when the acceptance rate is substantially higher.
@@ 4 Experiment
We conduct experiments in experimental and production-ready environments, using Vicuna 7B, 13B, 33B models as base LLMs. First, using PyTorch, we compare ReDrafter with state-of-the-art speculative decoding methods on an Nvidia H100 GPU in Section 4.1. Next, we validate ReDrafter’s performance gain in a production-ready environment on an inference server using TensorRT-LLM on H100, leveraging tensor parallelism and continuous batching in Section 4.2. Moreover, we investigate the on-device deployment of ReDrafter on Metal GPU using MLX, demonstrating its ability to accelerate on-device inference with limited computational resources in Section 4.3.
@@ 4.1 Pytorch Benchmark
MT-bench, temperature=1
Method Vicuna 7B Vicuna 13B Vicuna 33B
Speedup Tokens/Step Speedup Tokens/Step Speedup Tokens/Step
Medusa 2.33x 2.55 2.36x 2.61 2.52x 2.59
EAGLE 2.31x 3.51 2.27x 3.55 2.63x 3.48
ReDrafter 3.50x 5.31 3.51x 5.29 3.27x 4.69
AlpacaEval, temperature=0
@@
AlpacaEval, temperature=1
Method Vicuna 7B Vicuna 13B Vicuna 33B
Speedup Tokens/Step Speedup Tokens/Step Speedup Tokens/Step
Medusa 2.26x 2.42 2.24x 2.50 2.34x 2.43
EAGLE 2.20x 3.29 2.16x 3.34 2.41x 3.16
ReDrafter 3.53x 5.31 3.50x 5.22 3.32x 4.71
We compare ReDrafter with Medusa (Cai et al., 2024) and EAGLE (Li et al., 2024a) using PyTorch. For each method, we report speedups relative to auto-regressive decoding, as well as the average number of tokens accepted by LLM per generation step (Tokens/Step) on MT-Bench (Zheng et al., 2024) and Alpaca (Dubois et al., 2024). We conduct experiments at two different temperatures, 0 and 1. Temperature=0 represents the greedy approach (i.e., token matches), while temperature=1 corresponds to rejection sampling.
@@
Table 1 compares different methods. When temperature equals to 1, ReDrafter outperforms Medusa and EAGLE across all LLMs and evaluation datasets, achieving both the highest speedup and highest Tokens/Step. When temperature is 0, ReDrafter attains the highest speedup and Tokens/Step with Vicuna 7B and 13B. For 33B, ReDrafter achieves the highest Tokens/Step, though its speedup is slightly lower than EAGLE’s. Notably, ReDrafter achieves approximately 30% greater speedup at temperature 1 compared to temperature 0. This improved performance at temperature 1 can be attributed to the rejection sampling mechanism. At higher temperatures, the distributions of the LLM and draft model become smoother, increasing the likelihood of draft tokens being accepted by rejection sampling.
@@
4.2 TensorRT-LLM-based Cuda serving
Table 2: TensorRT-LLM experiments on MT-bench with continuous batching at varying request rates, using fixed settings of TP=1, beam width=6, beam length=5, and temperature=1.
Request Rate Queries Per Second (QPS)
Model QPS=0.1 QPS=0.5 QPS=1.0
ms/token Speedup ms/token Speedup ms/token Speedup
V 7B 2.41 2.09x 2.48 2.11x 2.76 1.57x
V 13B 3.41 2.31x 3.61 1.71x 5.43 1.22x
V 33B 5.93 2.56x 8.41 1.49x 20.04 1.37x
Table 3: TensorRT-LLM experiments on MT-bench using static batching with varying tensor parallelism (TP). We fix batch size=1, beam width=6, beam length=5, and temperature=1.
# Tensor Parallelism (TP) of LLM
Model TP=1 TP=2 TP=4
TPS Speedup TPS Speedup TPS Speedup
V 7B 323.12 2.19x 373.27 1.82x 403.87 1.68x
V 13B 198.64 2.30x 251.55 1.89x 312.79 1.84x
V 33B 97.50 2.51x 137.47 2.15x 171.11 1.93x
To integrate our invention into production-ready environments, we developed a TensoRT-LLM implementation to address high traffic and support long context. We demonstrate significantly improved the CUDA serving performance using Nvidia H100 GPUs.
In production, we run an inference service across multiple GPUs to handle concurrent user requests. To maintain service quality amid fluctuating traffic, latency and throughput are key metrics. Additionally, users typically focus on the latency of the first and last generated tokens. Previous works (Kwon et al., 2023; Kumar et al., 2024; Zhao et al., 2024) introduced techniques such as continuous batching to meet these demands. Given that TensorRT-LLM incorporates these methods, we use our own TensorRT-LLM implementation on 8 GPUs to empirically evaluate the performance improvements achieved with ReDrafter, demonstrated in Table 2.
We observe that ReDrafter significantly enhances the performance across all levels of Queries Per Second (QPS). At QPS 0.1, the speedup ranges from an impressive 2.09x to 2.56x, comparing with the highly optimized auto-regressive baseline in TensorRT-LLM. When QPS raise to 1, the speedup dropped to 1.22x to 1.57x. This is expected, since higher QPS leads to a larger dynamic batch size, pushing the system to a compute limit. We conduct an in-depth investigation of this phenomenon in an experimental environment in Section 4.4.1.
Another requirement for a production system is support for long context. To achieve this, serving systems often shard the model across multiple GPUs, even when it could fit in a single GPU’s DRAM, to free up memory for the context. In LLM serving, this sharding is typically done through tensor parallelism (TP, Shoeybi et al. (2019); Karakus et al. (2021); Ko et al. (2021); Gao et al. (2022)). We empirically examine the overhead associated with TP. Table 3 presents TensorRT-LLM Tokens Per Second and Speedup with varying TPs on MT-Bench. We disable continuous batching and set the batch size to 1 to eliminate traffic variability.
For TP=1, ReDrafter accelerates generation by 2.19x to 2.51x, while for TP=4, the speedup ranges from 1.68x to 1.93x. The speedup is less pronounced with higher TPs, when LLM forward pass becomes faster, while the RNN draft model is not fully optimized with TP. The overhead for the draft model constitutes a larger proportion of the overall wall-time, leading to less salient speedup.
Additionally, the observed speedup is not as significant as in the PyTorch experiments (3.27x to 3.50x) in Section 4.1 using the same checkpoint. This discrepancy arises because applying TP to the RNN draft model increases the number of all-gather operations when collecting intermediate tensors from multiple GPUs during beam search, incurring higher overhead compared to PyTorch. Additionally, the AR implementation in TensorRT-LLM is highly optimized, which limits the potential speedup achieved by ReDrafter.
@@ 4.3 MLX-based on device inference
Table 4: MLX experiments with varying beam width (BW). We fix temperature=0, batch size=1, beam length=4. See A.2 for more experimental setup details.
@@ A.2.2 Performance Tuning: MLX and Apple Silicon
Before starting the MLX implementation, we had experience with PyTorch and TensorRT-LLM. However, working on our first MLX project revealed that many lessons we learned from programming CUDA no longer apply to MLX and Apple Silicon. Below, we share some of the key lessons learned from this journey.

### v5

@@ Abstract
We present Recurrent Drafter (ReDrafter), an advanced speculative decoding approach that achieves state-of-the-art speedup for large language models (LLMs) inference. The performance gains are driven by three key aspects: (1) leveraging a recurrent neural network (RNN) as the draft model conditioning on LLM’s hidden states, (2) applying a dynamic tree attention algorithm over beam search results to eliminate duplicated prefixes in candidate sequences, and (3) training through knowledge distillation from the LLM. ReDrafter accelerates Vicuna inference in MT-Bench by up to 2.8x with a PyTorch implementation on Nvidia H100 GPUs. To demonstrate its practicality in real environments, we also validated its effectiveness for on-device applications by implementing the approach in MLX and benchmarking performance on Metal GPUs in Apple Silicon chips, achieving up to 2.3x speedup. We summarize our experimental results in Figure 1.
@@ 1 Introduction
Figure 1: Inference speedup on MT-bench for Vicuna models are shown for three ReDrafter implementations: (left) PyTorch on Nvidia H100 GPU, (right) and MLX on Apple’s M2 Ultra Metal GPU. We compare with EAGLE (Li et al., 2024a), Medusa (Cai et al., 2024), and auto-regressive in PyTorch implementations.
@@
In a PyTorch implementation, ReDrafter accelerates Vicuna inference by up to 2.8x compared to the autoregressive method in MT-Bench on Nvidia H100 GPUs, achieving state-of-the-art performance. Additionally, we validate the effectiveness of ReDrafter using MLX on Metal GPUs within Apple Silicon chips. Despite the limited compute resources in this setup, we observed a memory bottleneck. ReDrafter effectively mitigates this bottleneck, resulting in up to 2.3x speedup, demonstrating its capability to optimize performance in resource-constrained environments.
@@ 2 Related Work on Speculative Decoding
Another thread of approaches employs a unified strategy by attaching the draft model to the LLM, making them dependent (Stern et al., 2018; Santilli et al., 2023; Cai et al., 2024). This is an efficient design choice when the draft model is not intended for standalone use, allowing it to leverage additional computational resources for local prediction by conditioning on the LLM. Among these methods, Cai et al. (2024) proposes to use $T$ independent draft heads to predict next $T$ tokens, utilizing GPU’s extra parallel computing power. This excessive computational effort may not yield proportional speedups, as independent predictions become less accurate as $T$ increases, resulting in suboptimal predictive accuracy and a lower acceptance rate from the LLM. An alternative approach is to enhance prediction accuracy by incorporating recurrence into the draft model to capture local dependencies (Bhendawade et al., 2024; Li et al., 2024a; b; ankner2024hydra). However, the reduced parallelism due to recurrence leads to lower GPU utilization, introducing overhead that diminishes speedup gains, even when the acceptance rate is substantially higher.
@@ 4 Experiment
We conduct experiments in experimental and production-ready environments, using Vicuna 7B, 13B, 33B models as base LLMs. First, using PyTorch, we compare ReDrafter with state-of-the-art speculative decoding methods on an Nvidia H100 GPU in Section 4.1. Next, we investigate the on-device deployment of ReDrafter on Metal GPU using MLX, demonstrating its ability to accelerate on-device inference with limited computational resources in Section 4.2.
@@ 4.1 Pytorch Benchmark
AlpacaEval
@@
We compare ReDrafter with Medusa (Cai et al., 2024) and EAGLE (Li et al., 2024a) using PyTorch. For each method, we report speedups relative to auto-regressive decoding, as well as the average number of tokens accepted by LLM per generation step (Tokens/Step) on MT-Bench (Zheng et al., 2024) and Alpaca (Dubois et al., 2024). We mainly conduct our experiments with temperature$=0$, i.e., greedy decoding.
@@
Table 1 compares different methods. ReDrafter attains the highest speedup and Tokens/Step with Vicuna 7B and 13B. For 33B, ReDrafter achieves the highest Tokens/Step, though its speedup is slightly lower than EAGLE’s.
@@ 4.2 MLX-based on device inference
Table 2: MLX experiments with varying beam width (BW). We fix batch size=1, beam length=4. See A.2 for more experimental setup details.
@@ A.2.2 Performance Tuning: MLX and Apple Silicon
Before starting the MLX implementation, we had experience with PyTorch with CUDA. However, working on our first MLX project revealed that many lessons we learned from programming CUDA no longer apply to MLX and Apple Silicon. Below, we share some of the key lessons learned from this journey.